Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 228     Entry time: 06 Jun 2023, 23:11
Author: Loic Le Guyader 
Group: SCS 
Subject: late shift summary 

achievements

  • aligned the beam through up to the XOX imager
  • prepared FFT for reflectivity measurement
  • tested new DSSC gain files for Tb M5 edge

problems

  • in-vacuum piezo motor for the THz parabola stopped working, venting XOX required for troubleshooting, not using the X-ray beam this night

 

ELOG V3.1.4-7c3fd00